OpenStudy (anonymous):

y = -1/3 x + 2 and x + 3y = 3

OpenStudy (anonymous):

Is t -1/(3x)?

jhonyy9 (jhonyy9):

so than 3y=-x+6 and x+3y=3 so 3y+x=6 3y+x=3 -------- subtract (3y-3y)+(x-x)=(6-3) 0 + 0 = 3 0 = 3 havent solution